### v1

# An EPTAS for multiprocessor scheduling with rejection under a machine cost constraint

Abstract
We study the multiprocessor scheduling with rejection problem under a machine cost constraint. In this problem, each job is either rejected with a rejection penalty or; accepted and scheduled on one of the machines for processing. The machine cost is proportional to the total processing time of the jobs scheduled on it. The problem aims to minimize the makespan of accepted jobs plus the total rejection penalty of rejected jobs while the total machine cost does not exceed a given upper bound. We present a simple $2$-approximation algorithm for the problem and we achieve an EPTAS when the number $m$ of machines is a fixed constant.
Keywords: Scheduling within rejection; rejection penalty; green manufacturing; approximation scheme
1 Introduction
In this paper, we consider the multiprocessor scheduling with rejection problem under a machine cost constraint [12]. Formally, we are given a set of jobs $J=\{J_{1},\ldots,J_{n}\}$ and a set of identical machines $M=\{M_{1},\ldots,M_{m}\}$. A job $J_{j}$ is either rejected with a rejection penalty $e_{j}$ or accepted and scheduled on some machine with a processing time $p_{j}$. Each machine $M_{i}$ is associated with an energy cost rate $r_{i}$ and the machine cost of $M_{i}$ is exactly $r_{i}$ times the maximum completion time of the jobs scheduled on $M_{i}$. The total machine cost is limited and cannot exceed a given upper bound $\hat{U}$. The problem aims to find a solution to minimize the makespan of accepted jobs plus the total rejection penalty of rejected jobs while the machine cost constraint is satisfied. The problem is denoted as $Pm\mid rej,\hat{U},r_{i}\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ where $m$ is a fixed constant; $r_{i}$ can be non-uniform; $A$ and $R$ are the accepted and rejected jobs, respectively. If $m$ is part of the input, then we use $P$ to replace $Pm$ and if $r_{i}$ is uniform, then we use $r_{i}=r$ to replace $r_{i}$ in the above three-field notation.
The problem is motivated by the applications in the green manufacturing industry [12], which recently attracts much attention because of the growth concerns about environmental challenges. Generally speaking, green manufacturing industry aims to maximize the resource utilities while reduces the environmental damage as much as possible [14]. The machine cost is an important indicator since it may measure the energy cost for the machines, for example, the clean-up cost of pollutants caused by the machines [14]. The authors in [14] initiated the research for scheduling problems in the green manufacturing industry, i.e., $P\mid\hat{U},r_{i}\mid C_{\max}$ and $P\mid\hat{U},r_{i}\mid\sum C_{j}$ where $C_{\max}$ stands for the makespan objective and $\sum C_{j}$ stands for the min-sum objective. Recently, Kong et al. [12] generalized the problem to $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ by assuming jobs can be either rejected or accepted; but the machine energy cost rate is uniform.
For a minimization problem, an algorithm $A$ is $\alpha$-approximation if for every instance $I$ of the problem, it runs in polynomial time in the input size of $I$, and always computes a feasible solution whose objective value is at most $\alpha$ times the optimal value of $I$. A polynomial-time approximation scheme (PTAS) is a family of approximation algorithms such that, for any $\epsilon>0$, one of which is $(1+\epsilon)$-approximation. A PTAS is an efficient PTAS (EPTAS) if the running time of each algorithm in the family can be expressed as $f(\frac{1}{\epsilon})n^{O(1)}$ where $f$ is a polynomial-time computable function. An EPTAS becomes a fully PTAS (FPTAS) if $f(\frac{1}{\epsilon})=(\frac{1}{\epsilon})^{O(1)}$.
To our knowledge, the only current best $2$-approximation algorithm for $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ is due to reference [12]. We next review the related works of $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$. Note that when $\hat{U}=\infty$ and each rejection penalty $e_{j}=\infty$, the scheduler is forced to accept all jobs and thus the problem is reduced to the classic multiprocessor scheduling problem $P\mid\mid C_{\max}$ [6]. The problem $P\mid\mid C_{\max}$ is one of the most significant NP-hard problems that has been well studied in the literature. For any fixed $m\geq 2$, $Pm\mid\mid C_{\max}$ is weakly NP-hard [6], for which Sahni presented an FPTAS [15]. Moreover, $P\mid\mid C_{\max}$ has been shown to be strongly NP-hard [11]. Graham [9] designed the first algorithm, called list scheduling, for $P\mid\mid C_{\max}$ in 1966, which always schedules a job on a machine with the least load and achieves $(2-\frac{1}{m})$-approximation. Afterwards, the same author [10] showed that if the jobs are sorted in the non-increasing order of their processing times, the approximation ratio can be improved to $\frac{4}{3}-\frac{1}{3m}$. Finally, Hochbaum and Shmoys [11] resolved this problem by presenting a PTAS.
When $\hat{U}=\infty$, the problem $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ is reduced to the scheduling with rejection problem proposed by Bartal et al. [3]. The authors in [3] presented an FPTAS when $m$ is a fixed constant and a PTAS when $m$ is part of the input. The authors in [12] initiated the research on a general $\hat{U}$ and presented a $2$-approximation algorithm for both a single machine and multiple machines. The interested readers may refer to the survey [16] for more related problems of scheduling with rejection and their corresponding results.
1.1 Other related works
Note that in a scheduling with rejection problem, the rejection penalty is directly added to the objective if a job is rejected. Many different settings are also considered in the literature. For example, in a scheduling with testing problem, each job $J_{j}$ is given an upper bound $u_{j}$ on the exact processing time $p_{j}$ and a testing time $t_{j}$. A job $J_{j}$ is either untested and processed for $u_{j}$ time or; tested for $t_{j}$ time and processed for $p_{j}$ time on some machine. It is worth mentioning that the testing time $t_{j}$ is added to the processing time of $J_{j}$ instead of the objectives. Dürr et al. [5] first introduced the scheduling with testing problem on a single machine where the testing times are uniform and the goal is to minimize the total completion time. Albers and Eckl [1] extended the model by allowing the testing times to be non-uniform. Moreover, they [2] also pioneered research on minimizing the makespan on multiple machines. Other algorithmic results for the scheduling with testing problems can be found in the references [8, 7].
The scheduling with compression problem is another variant in which the testing is treated as a monetary cost. In other words, the processing time of a job $J_{j}$ can be compressed to a certain value $p_{j,k}$ when a cost $c_{k}$ is paid from a given budget $B$. In [17], Zhang and Zhang achieved an FPTAS for scheduling with compression with the makespan objective when the number of machines is a fixed number and each job has a constant number of compressible processing times. When every job allows for two processing times $p_{j,1}\geq p_{j,2}$ and $c_{j,2}=0$, the problem is reduced to the scheduling within budget problem, studied by Damerius et al. [4]. Damerius et al. [4] achieved a novel PTAS for the min-sum objective on a single machine.
1.2 Our results
Our main contribution is an EPTAS for the problem $Pm\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$, where the number of machines is a fixed constant. Before the EPTAS, we present a new and simpler $2$-approximation algorithm than the one in [12] for $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ by guessing the maximum processing time of accepted jobs and the maximum rejection penalty of rejected jobs. Our EPTAS first distinguishes the risky jobs whose rejection penalties are relatively big and constructs another instance such that we can easily make the rejection decisions for risky jobs. We next carefully determine the assignments of the ”large” risky jobs that are accepted and present a linear program to deal with the remaining jobs. The jobs with a fractional solution are either rejected or; accepted and appended to the schedule, leading to a small increase on the cost.
The rest of the paper is organized as follows. In Section 2, we give the formal definitions and notation for the problem. In Section 3, we present a new and simple $2$-approximation algorithm and then in Section 4, the EPTAS for $Pm\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$. Finally, we conclude our paper and introduce some other research problems that can be studied in the future.
2 Preliminaries
We first present the formal definitions for the scheduling with rejection problem. An instance $I$ of the problem consists of a set of $n$ jobs $J=\{J_{1},\ldots,J_{n}\}$ and a set of $m$ identical machines $M=\{M_{1},\ldots,M_{m}\}$. Each job $J_{j}$ has a known processing time $p_{j}$ and a rejection penalty $e_{j}$. The scheduler needs to make a rejection decision for every job $J_{j}$ and in other words, a job $J_{j}$ is either rejected or accepted. If $J_{j}$ is rejected, then the rejection penalty should be paid; otherwise, $J_{j}$ is accepted and should be scheduled on some machine for processing. The completion time of a machine $M_{i}$, denoted as $t_{i}$, is the maximum completion time of the jobs scheduled on $M_{i}$, that is, the total processing time of the jobs scheduled on $M_{i}$. The makespan of a schedule is the maximum completion time of all machines.
In the green manufacturing industry, the energy cost caused by a machine is usually proportional to its completion time and the total energy cost cannot exceed a given threshold $\hat{U}$. We assume that the energy cost rate for each machine is uniform and is denoted as $r$. Therefore, we have the constraint ”$r\sum_{i=1}^{m}t_{i}\leq\hat{U}$” on the total energy cost [12]. To simplify the notations, we let $U=\hat{U}/r$ and the constraint becomes $\sum_{i=1}^{m}t_{i}\leq U$, which is equivalent to saying that the total processing time of accepted jobs is bounded by $U$. A feasible solution needs to determine the rejection decision for each job and schedule the accepted jobs on the machines while the constraint is satisfied. We define the cost $Z$ of a feasible solution is the makespan of accepted jobs plus the total rejection penalty of rejected jobs. Our goal is to find a feasible solution such that the cost is minimized. If we let $A$ and $R$ denote the accepted and rejected jobs, respectively, then the problem is denoted as $Pm\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ when the number $m$ of machines is a fixed constant or $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ when $m$ is part of the input.
Given an instance $I$ of $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$, we fix an optimal solution of $I$ for discussion and use $A^{*}(I)$ and $R^{*}(I)$ to denote the accepted and rejected job set in the optimal solution, respectively. Moreover, we let $C^{*}_{\max}(I)$ denote the makespan of the optimal solution and $Z^{*}(I)$ denote the optimal cost. Similarly, for a specific solution produced by an algorithm, we let $C_{\max}(I)$ denote the makespan and $Z(I)$ denote the cost of the solution. We drop “$(I)$” from these notations if the instance $I$ is clear from the context.
For the classic multiprocessor scheduling problem $P\mid\mid C_{\max}$, the list scheduling rule always assigns a job to the least-loaded machine and is a $2$-approximation. We employ this rule in the $2$-approximation algorithm for $P\mid rej,U,r_{i}=r\mid C_{\max}+\sum_{J_{j}\in R}e_{j}$ in Section 3
Definition 1.
(List scheduling rule [9]) The List scheduling rule, LS rule for short, always schedules an unassigned job to a machine with the least load.
3 A new $2$-approximation for $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$
We remark that the algorithm in this section can be applied to $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ where the number $m$ of machines is part of the input. Recall that the accepted and rejected job sets are denoted as $A^{*},R^{*}$ in the optimal solution, respectively. Let $p^{*}_{\max}$ denote the maximum processing time of the jobs in $A^{*}$ and $e^{*}_{\max}$ denote the maximum rejection penalty of the jobs in $R^{*}$, respectively. The key idea is to guess the values of $p^{*}_{\max}$ and $e^{*}_{\max}$ so that we can immediately reject the jobs with $p_{j}>p^{*}_{\max}$ and accept the jobs with $e_{j}>e^{*}_{\max}$. The following definition is directly from the fact $p^{*}_{\max}\in\{p_{1},\ldots,p_{n}\}$ and $e^{*}_{\max}\in\{e_{1},\ldots,e_{n}\}$.
Definition 2.
A tuple $(p,e)$ is a guess if $p\in\{p_{1},\ldots,p_{n}\}$ and $e\in\{e_{1},\ldots,e_{n}\}$.
It is easy to see that the number of distinct guesses is bounded by $n^{2}$. Note that we do not know $p^{*}_{\max}$ and $e^{*}_{\max}$ and we choose to enumerate all possible guesses. In the following, we use a specific guess $(p_{\ell},e_{k})$ for further discussion. Let
$A_{1}=\{J_{j}\in J:e_{j}>e_{k}\}\mbox{ and }R_{1}=\{J_{j}\notin A_{1}:p_{j}>p_{\ell}\mbox{ or }p_{j}>me_{j}\}.$ (1)
We accept all the jobs in $A_{1}$ and reject all the jobs in $R_{1}$. If the guess is correct, i.e., $(p_{\ell},e_{k})=(p^{*}_{\max},e^{*}_{\max})$, then the jobs in $A_{1}$ are accepted and the jobs in $R_{1}$ with $p_{j}>p_{\ell}$ are rejected in the optimal solution. Therefore, a necessary condition for a correct guess is $U>\sum_{J_{j}\in A_{1}}p_{j}$, which induces the following definition.
Definition 3.
A guess is valid if its corresponding set $A_{1}$ satisfies $U>\sum_{J_{j}\in A_{1}}p_{j}$.
If $(p_{\ell},e_{k})$ is not a valid guess, then we skip it and try another guess. Recall that we accept all the jobs in $A_{1}$ and reject all the jobs in $R_{1}$ with $p_{j}>p_{\ell}$, the same as in the optimal solution if the guess is correct. Moreover, we also reject the jobs with $p_{j}>me_{j}$ because the rejection penalty is relatively smaller than the processing time.
For ease of presentation, we let $X=J\setminus(A_{1}\cup R_{1})$. We next present a linear program to determine the rejection decisions for most jobs in $X$. In more detail, let $x_{j}\in\{0,1\}$ denote a binary decision variable for the job $J_{j}\in X$ such that $x_{j}=1$ if and only if $J_{j}$ is rejected by the algorithm. Moreover, let $y_{j}=1-x_{j}$, where $y_{j}$ is also a binary decision variable such that $y_{j}=1$ indicates that $J_{j}$ is accepted. The linear program is presented as follows:
$\min\quad\frac{1}{m}$ $\sum_{J_{j}\in X}y_{j}p_{j}+\sum_{J_{j}\in X}x_{j}e_{j}$ (2)
$\sum_{J_{j}\in X}y_{j}p_{j}\leq U-\sum_{J_{j}\in A_{1}}p_{j},$
$x_{j}+y_{j}=1,\quad\forall J_{j}\in X.$
We remark that the objective function is the average processing time on the machines plus the total rejection penalty, which is close to the problem objective $C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$. The first constraint of LP(2) is because the total processing time of the accepted jobs cannot exceed the upper bound $U$ and the jobs in $A_{1}$ have been accepted. Since the guess $(p_{\ell},e_{k})$ is valid, by Definition 3, the linear program has at least one feasible solution, i.e., $x_{j}=1$ for every job $J_{j}\in X$. Note that there are $2|X|\leq 2n$ variables and $|X|+1\leq n+1$ constraints. The linear program is polynomial-time solvable in $n$ because the time complexity of a linear program is polynomial in the number of continuous variables and polynomial in the number of constraints [13].
We consider a basic solution of LP(2). We partition the jobs in $X$ into three sets based on the decision variables in the basic solution:
$X_{1}=\{J_{j}\in X:y_{j}=1\},X_{2}=\{J_{j}\in X:x_{j}=1\}\mbox{ and }X_{3}=X\setminus(X_{1}\cup X_{2}).$ (3)
Note that $\sum_{J_{j}\in X_{1}}p_{j}\leq U-\sum_{J_{j}\in A_{1}}p_{j}$. We reject all the jobs in $X_{2}\cup X_{3}$, in other words, the rejected job set is exactly $R_{1}\cup X_{2}\cup X_{3}$. Finally, we apply the LS rule in Definition 1 to schedule the accepted jobs. The pseudocode of our algorithm is presented as follows.
Algorithm 1 Algorithm Approx1 for $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$
Input: An instance $J=\{J_{1},\ldots,J_{n}\}$ and $m$ machines;
for Each guess $(p_{\ell},e_{k})$ do
     Define two sets $A_{1},R_{1}$ as in Eq.(1) and $X=J\setminus(A_{1}\cup R_{1})$.
     if $(p_{\ell},e_{k})$ is valid then
         Construct the linear program as in LP(2) and compute a basic solution.
         Partition the jobs in $X$ into three sets $X_{1},X_{2},X_{3}$ as in Eq.(3).
         Reject all the jobs in $X_{2}\cup X_{3}\cup R_{1}$ and apply LS rule to schedule all accepted jobs.
     end if
end for
Return the solution with minimum cost corresponding to the valid guesses.
We next analyze the approximation ratio and without loss of generality, we assume the guess is correct. The following lemma shows that at most one job belongs to $X_{3}$.
Lemma 4.
$|X_{3}|\leq 1$.
Proof.
Since the number of constraints of LP(2) is $|X|+1$, there are at most $|X|+1$ positive values in the basic solution. By Eq.(3), we have exactly $|X_{1}|+|X_{2}|+2|X_{3}|=|X|+|X_{3}|$ positive values. It follows from $|X|+|X_{3}|\leq|X|+1$ that $|X_{3}|\leq 1$. The proof is done. ∎
Lemma 5.
If the guess is correct, then we have
$\frac{1}{m}\sum_{J_{j}\in X_{1}}p_{j}+\sum_{J_{j}\in X_{2}}e_{j}\leq\frac{1}{m}\sum_{J_{j}\in A^{*}\cap X}p_{j}+\sum_{J_{j}\in R^{*}\cap X}e_{j}$ (4)
Proof.
Consider an assignment of $x_{j}$ for each job $J_{j}\in X$: $y_{j}=1$ if and only if $J_{j}\in A^{*}\cap X$. Once the value of $y_{j}$ is fixed, we set $x_{j}=1-y_{j}$. Note that such an assignment is a feasible solution of LP(2) if the guess is correct. The objective value of the assignment is exactly the right side of Eq.(4), which is an upper bound on the optimal value of LP(2). The proof is done since the left side of Eq.(4) is not greater than the optimal value of LP(2). ∎
Theorem 6.
Algorithm Approx1 is a $2$-approximation algorithm for $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$.
Proof.
Since the jobs in $A^{*}$ are accepted, the makespan is at least $\max\{p^{*}_{\max},\frac{1}{m}\sum_{J_{j}\in A^{*}}p_{j}\}$. Therefore, we have
$Z^{*}\geq\max\{p^{*}_{\max},\frac{1}{m}\sum_{J_{j}\in A^{*}}p_{j}\}+\sum_{J_{j}\in R^{*}}e_{j}.$ (5)
Suppose the guess is correct and the rejection penalty is clearly $\sum_{J_{j}\in R_{1}\cup X_{2}\cup X_{3}}e_{j}$. By Lemma 4, $|X_{3}|\leq 1$ and since $X_{3}\subseteq X$, we have $\sum_{J_{j}\in X_{3}}e_{j}\leq e^{*}_{\max}$. Recall that any job $J_{j}$ in $R_{1}$ with $p_{j}>p^{*}_{\max}$ is also rejected in the optimal solution. It follows that for any job $J_{j}\in R_{1}\cap A^{*}$, we have $p_{j}>me_{j}$. Therefore, the rejection penalty is at most
$\sum_{J_{j}\in R_{1}\cup X_{2}\cup X_{3}}e_{j}$ $=$ $\sum_{J_{j}\in R_{1}\cap A^{*}}e_{j}+\sum_{J_{j}\in R_{1}\cap R^{*}}e_{j}+\sum_{J_{j}\in X_{2}\cup X_{3}}e_{j}$ (6)
$\leq$ $\frac{1}{m}\sum_{J_{j}\in R_{1}\cap A^{*}}p_{j}+\sum_{J_{j}\in R_{1}\cap R^{*}}e_{j}+\sum_{J_{j}\in X_{2}}e_{j}+e^{*}_{\max}.$
We next analyze the makespan of the schedule, which we assume, is determined by the job $J_{t}$. Since $J_{t}\notin R_{1}$, $p_{t}\leq p^{*}_{\max}$. Recall that $J_{t}$ is scheduled on a least-loaded machine. Therefore, we have
$C_{\max}\leq\frac{1}{m}\sum_{J_{j}\in A_{1}\cup X_{1}}p_{j}+p_{t}\leq\frac{1}{m}\sum_{J_{j}\in A_{1}\cup X_{1}}p_{j}+p^{*}_{\max}.$ (7)
Using Eqs.(5)-(7) and Lemma 5, the cost is at most
$Z$ $=C_{\max}+\sum_{J_{j}\in R_{1}\cup X_{2}\cup X_{3}}e_{j}$
$\leq\frac{1}{m}\sum_{J_{j}\in R_{1}\cap A^{*}}p_{j}+\sum_{J_{j}\in R_{1}\cap R^{*}}e_{j}+\sum_{J_{j}\in X_{2}}e_{j}+e^{*}_{\max}+\frac{1}{m}\sum_{J_{j}\in A_{1}\cup X_{1}}p_{j}+p^{*}_{\max}$ (By Eqs.(6, 7))
$=\frac{1}{m}\sum_{J_{j}\in R_{1}\cap A^{*}}p_{j}+\frac{1}{m}\sum_{J_{j}\in A_{1}}p_{j}+\sum_{J_{j}\in R_{1}\cap R^{*}}e_{j}+e^{*}_{\max}+p^{*}_{\max}+\frac{1}{m}\sum_{J_{j}\in X_{1}}p_{j}+\sum_{J_{j}\in X_{2}}e_{j}$
$\leq\frac{1}{m}\sum_{J_{j}\in R_{1}\cap A^{*}}p_{j}+\frac{1}{m}\sum_{J_{j}\in A_{1}}p_{j}+\sum_{J_{j}\in R_{1}\cap R^{*}}e_{j}+e^{*}_{\max}+p^{*}_{\max}+\frac{1}{m}\sum_{J_{j}\in A^{*}\cap X}p_{j}+\sum_{J_{j}\in R^{*}\cap X}e_{j}$ (By Lemma 5)
$=\frac{1}{m}\sum_{J_{j}\in A^{*}}p_{j}+p^{*}_{\max}+\sum_{J_{j}\in R^{*}}e_{j}+e^{*}_{\max}$
$\leq 2Z^{*}.$ (By Eq.(5))
The proof is finished. ∎
4 An EPTAS for $Pm\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$
We fix an $10<\epsilon<$ 1 where $\frac{1}{\epsilon}$ is an integer. We assume $m$ is a fixed constant and present an EPTAS for $Pm\mid rej,U,r_{i}=r\mid C_{\max}+\sum_{J_{j}\in R}e_{j}$.
4.1 Data normalization
We fix an instance $I$ of $Pm\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ for discussion. The first step in the EPTAS is to call Approx1 and by Theorem 6, it produces a feasible solution such that the cost $Z(I)$ satisfies $\frac{1}{2}Z(I)\leq Z^{*}(I)\leq Z(I)$. We normalize the instance by dividing each $p_{j},e_{j}$ by the quantity $Z(I)$, and thus without loss of generality, we assume
$\frac{1}{2}\leq Z^{*}(I)\leq 1=Z(I).$ (8)
We set up a fixed constant $0<\Delta<\epsilon$ in Eq.(9) based on the values of $m,\epsilon$. Since $\frac{1}{\epsilon}$ is an integer, $\frac{1}{\Delta}$ is an integer as well.
$\Delta=\frac{\epsilon}{4m+12}.$ (9)
We partition the jobs into risky and safe using the constant $\Delta$.
Definition 7.
A job $J_{j}$ is risky if $e_{j}\geq\Delta$ and otherwise, $J_{j}$ is safe.
Note that if $e_{j}<\Delta$, then the rejection penalty is relatively small and thus it is safe to reject it.
Lemma 8.
At most $\frac{1}{\Delta}$ risky jobs are rejected in the optimal solution.
Proof.
The proof is trivial since otherwise, the optimal cost is greater than $1$, which contradicts Eq.(8). ∎
4.2 Restricted instance
In this section, we construct another instance $I^{\prime}$ from the input instance $I$ but the number of distinct rejection penalty becomes a fixed constant. To achieve the goal, we set up another fixed constant $0<\delta<\Delta<\epsilon$ in Eq.(10). We note that $\frac{1}{\delta}$ is an integer since $\frac{1}{\Delta}$ and $\frac{1}{\epsilon}$ are both integers.
$\delta=\frac{\epsilon\Delta}{2}=\frac{\epsilon^{2}}{8m+24}.$ (10)
Definition 9.
(Restricted instance $I^{\prime}$) The restricted instance $I^{\prime}$ is constructed from the input instance $I$ by rounding the rejection penalty of each risky job up to the least multiple of $\delta$.
We let $J^{\prime}_{j}$ be the job in $I^{\prime}$ corresponding to the job $J_{j}$ in $I$, whose processing time and rejection penalty are denoted as $p^{\prime}_{j}$ and $e^{\prime}_{j}$, respectively. By the construction of $I^{\prime}$, we have $p^{\prime}_{j}=p_{j}$ for every job $J_{j}$. If $J_{j}$ is safe, then $e^{\prime}_{j}=e_{j}$. Otherwise, $e_{j}\geq\Delta$ and thus, by Eq.(10), $e^{\prime}_{j}\leq e_{j}+\delta\leq(1+\frac{\epsilon}{2})e_{j}$ We conclude that for every job $J_{j}$, we have $e^{\prime}_{j}\leq(1+\frac{\epsilon}{2})e_{j}$. Moreover, we know $e^{\prime}_{j}\geq\Delta$ if and only if $e_{j}\geq\Delta$. In other words, $J^{\prime}_{j}$ is risky if and only if $J_{j}$ is risky. We conclude all those facts in the following remark.
Remark 10.
For the instances $I$ and $I^{\prime}$, we have the following three facts:
  • 1.
    $p^{\prime}_{j}=p_{j}$ for every job $J_{j}$;
  • 2.
    $e^{\prime}_{j}\leq(1+\frac{\epsilon}{2})e_{j}$ for every job $J_{j}$;
  • 3.
    $J_{j}$ is risky if and only if $J^{\prime}_{j}$ is risky.
Lemma 11.
There exists a feasible solution $\pi$ for the restricted instance $I^{\prime}$ such that
  • 1.
    At most $\frac{1}{\Delta}$ risky jobs are rejected;
  • 2.
    the cost is at most $(1+\frac{\epsilon}{2})Z^{*}(I)$.
Proof.
Consider the optimal solution $\pi^{*}$ for the instance $I$ and let $R^{*}$ denote the rejected job set. We construct $\pi$ from $\pi^{*}$ as follows: $J^{\prime}_{j}$ is rejected in $\pi$ if and only if $J_{j}$ is rejected in $\pi^{*}$. Once a job is accepted, then it will be scheduled on the same machine as in $\pi^{*}$. Clearly, $\pi$ is a feasible solution for $I^{\prime}$ since the processing times and the upper bound $U$ are not changed.
Note that we make the same rejection decision for each job as in $\pi^{*}$. By Lemma 8 and Statement 3 in Remark 10, at most $\frac{1}{\Delta}$ risky jobs are rejected and the first statement holds.
We next prove the second statement. By Statement 1 of Remark 10 and the construction of $\pi$, the makespan of $\pi$ is the same as $\pi^{*}$ and it is sufficient to show the total rejection penalty is bounded by $\sum_{J_{j}\in R^{*}}(1+\frac{\epsilon}{2})e_{j}$, which follows from Statement 2 in Remark 10. ∎
Note that in the restricted instance $I^{\prime}$, the rejection penalty of each risky job is a multiple of $\delta$ and is bounded by $1$. Therefore, we can partition the risky jobs into at most $\frac{1}{\delta}$ groups in the following definition.
Definition 12.
Let $S$ and $B$ denote the set of safe and risky jobs in the restricted instance $I^{\prime}$. We further partition the jobs in $B$ into $\frac{1}{\delta}$ subsets:
$B_{i}=\{J_{j}\in B:e_{j}=i\delta\},\mbox{ for every $i=1,2,\ldots,1/\delta$.}$
Since the jobs in $B_{i}$ have identical rejection penalties, we sort the jobs in $B_{i}$ by the non-increasing order of their processing times. This way, we have $J=S\cup\bigcup_{i=1}^{1/\delta}B_{i}$.
By Lemma 11 and Definition 12, at most $\frac{1}{\Delta}$ jobs in $\bigcup_{i=1}^{1/\delta}B_{i}$ are rejected.
Definition 13.
(Rejection types) A rejection type is a $\frac{1}{\delta}$-tuple $(k_{1},k_{2},\ldots,k_{1/\delta})$ such that each $k_{i}$ is a non-negative integer and $\sum_{i=1}^{1/\delta}k_{i}\leq 1/\Delta$.
A feasible solution of $I^{\prime}$ is associated with a rejection type $(k_{1},k_{2},\ldots,k_{1/\delta})$ if in the solution, exactly $k_{i}$ jobs of $B_{i}$ are rejected, for every $i=1,\ldots,1/\delta$.
Lemma 14.
The number of distinct rejection types is at most $N_{r}=(1/\Delta+1)^{1/\delta}$, which is a fixed constant when $m$ is a constant.
Proof.
Since $\sum_{i=1}^{1/\delta}k_{i}\leq 1/\Delta$ and $k_{i}$ is a non-negative integer, we have $k_{i}\in\{0,1,\ldots,1/\Delta\}$ for every $i=1,\ldots,1/\delta$. The lemma is proved because $1/\delta$ and $1/\Delta$ are fixed integers. ∎
Note that the feasible solution $\pi$ stated in Lemma 11 must be associated with a unique rejection type.
Lemma 15.
Suppose that $\pi$ is associated with the rejection type $(k_{1},k_{2},\ldots,k_{1/\delta})$. There exists a feasible solution $\pi^{\prime}$ for the restricted instance $I^{\prime}$ such that
  • 1.
    $\pi^{\prime}$ is associated with $(k_{1},k_{2},\ldots,k_{1/\delta})$, the same as $\pi$;
  • 2.
    the cost is at most $(1+\frac{\epsilon}{2})Z^{*}(I)$;
  • 3.
    the first $k_{i}$ jobs of $B_{i}$ are rejected for every $i=1,\ldots,1/\delta$.
Proof.
If the solution $\pi$ satisfies the third item, then by Lemma 11, $\pi$ is the desired solution.
Assume otherwise, in the solution $\pi$, there exists a pair of jobs $(p_{j},p_{t})$ in some job set $B_{i}$ such that $j\leq k_{i}\leq t$ and; $J_{j}$ is accepted and $J_{t}$ is rejected in $\pi$. By Definition 12, we have $e_{j}=e_{t}$ and $p_{j}\geq p_{t}$. We construct another solution $\pi_{1}$ from $\pi$ as follows: We accept $J_{t}$ and schedule it on the machine, where $J_{j}$ is assigned in $\pi$. We reject $J_{j}$ and for every job except $J_{j},J_{t}$, we make the same rejection decision and schedule it on the same machine as in $\pi$ if it is accepted. Clearly, the first item is satisfied.
Note that $p_{j}\geq p_{t}$ and thus $\pi_{1}$ is a feasible solution and by the construction of $\pi_{1}$, the makespan of $\pi_{1}$ is not greater than $\pi$. Since $e_{j}=e_{t}$, the rejection penalties of $\pi$ and $\pi_{1}$ are exactly the same. Therefore, the cost of $\pi_{1}$ does not exceed the cost of $\pi$. By Lemma 11, the second item in the lemma is satisfied. We can continue to ”swap” the rejection decisions of two jobs until the third item is satisfied, which leads to a desired solution. ∎
By Lemma 14, there are at most $N_{r}=(1/\Delta+1)^{1/\delta}$ distinct rejection types. Therefore, we are able to enumerate all rejection types to find the one associated with $\pi^{\prime}$ in Lemma 15. Given the rejection type $(k_{1},k_{2},\ldots,k_{1/\delta})$ of $\pi^{\prime}$, it is easy to determine the rejection decisions for all risky jobs. In more details, we reject the first $k_{i}$ jobs of $B_{i}$, for every $i=1,\ldots,1/\delta$ and accept the other risky jobs. We use the notations $A^{\prime}_{1},R^{\prime}_{1}$ to denote the accepted and rejected risky jobs, respectively.
Definition 16.
A job $J^{\prime}_{j}\in A^{\prime}_{1}$ is large if $p^{\prime}_{j}\geq\Delta$; otherwise, $J^{\prime}_{j}$ is tiny. Let $L$ and $T$ be the large and tiny job set and thus $A^{\prime}_{1}=L\cup T$.
Lemma 17.
There are at most $(1+\frac{\epsilon}{2})\frac{m}{\Delta}$ large jobs, i.e., $|L|\leq(1+\frac{\epsilon}{2})\frac{m}{\Delta}$.
Proof.
By Eq.(9), we have $(1+\frac{\epsilon}{2})\frac{1}{\Delta}=\frac{1}{\Delta}+2m+6$, which is an integer. Assume there are at least $(1+\frac{\epsilon}{2})\frac{m}{\Delta}+1$ large jobs and thus in the schedule $\pi^{\prime}$, there exists a machine where at least $(1+\frac{\epsilon}{2})\frac{1}{\Delta}+1$ large jobs are scheduled in $\pi^{\prime}$. By Eq.(8), the makespan of $\pi^{\prime}$ is at least $((1+\frac{\epsilon}{2})\frac{1}{\Delta}+1)\Delta>1+\frac{\epsilon}{2}\geq(1+\frac{\epsilon}{2})Z^{*}(I)$, which contradicts Lemma 15. ∎
Note that each large job must be assigned to some machine in $\pi^{\prime}$. The following definition describes all possible assignments for large jobs.
Definition 18.
(Large job assignment types) A large job assignment type is an $|L|$-tuple $(\ell_{1},\ldots,\ell_{|L|})$ such that each $\ell_{i}\in\{1,2,\ldots,m\}$, for every $i=1,\ldots,|L|$.
A schedule is associated with a large job assignment type $(\ell_{1},\ldots,\ell_{|L|})$ if the $i$-th large job is assigned to the machine $M_{\ell_{i}}$ for every $i=1,\ldots,|L|$.
Lemma 19.
The number of distinct large job assignment types is at most $m^{(1+\frac{\epsilon}{2})\frac{m}{\Delta}}$, which is a fixed constant when $m$ is a constant.
Proof.
Similarly to Lemma 14, the lemma follows from $\ell_{i}\in\{1,2,\ldots,m\}$ and $|L|\leq(1+\frac{\epsilon}{2})\frac{m}{\Delta}$. ∎
Clearly, the schedule $\pi^{\prime}$ is associated with a unique large job assignment and by Lemma 19, we are able to enumerate all possible large job assignments to find the one associated with $\pi^{\prime}$. At this time, we have determined the rejection decision for each risky job and also determined the schedule of each large job.
4.3 Schedule tiny and safe jobs
We remind the readers that all tiny jobs are accepted by our algorithm. In this section, we present a linear program to determine the schedule of most tiny jobs and; determine the rejection decisions of most safe jobs and the designated machine for each accepted safe job. For every safe job $J_{j}\in S$, let $x_{j}\in\{0,1\}$ be a binary decision variable such that $x_{j}=1$ if and only if $J_{j}$ is rejected in the solution $\pi^{\prime}$. Moreover, for every (tiny or safe) job $J_{j}\in T\cup S$ and every machine $M_{i}$, let $y_{ji}\in\{0,1\}$ be a binary decision variable, which is $1$ if and only if $J_{j}$ is scheduled on the machine $M_{i}$ in $\pi^{\prime}$. Recall that we schedule all large jobs according to a given large job assignment type and we let $t_{i}$ be the total processing time of large jobs assigned to the machine $M_{i}$, for every $i=1,\ldots,m$. We are now ready to present our linear program.
$\min$ $C_{\max}+\sum_{J_{j}\in S}x_{j}e_{j}$ (11)
$\sum_{i=1}^{m}y_{ji}=1$ $\forall J_{j}\in T$
$x_{j}+\sum_{i=1}^{m}y_{ji}=1$ $\forall J_{j}\in S$
$C_{\max}\geq t_{i}+\sum_{J_{j}\in T\cup S}y_{ji}p^{\prime}_{j}$ $\forall M_{i}$
$\sum_{J_{j}\in S}(1-x_{j})p_{j}\leq U-\sum_{J_{j}\in A^{\prime}_{1}}p^{\prime}_{j}.$
We next interpret the LP(11) in details. The first constraint indicates that every tiny job should be scheduled on some machine. Note that if a safe job $J_{j}$ is rejected, then $x_{j}=1$ and $y_{ji}=0$ for every machine $M_{i}$. Otherwise, $J_{j}$ is accepted and should be scheduled on some machine. Therefore, $x_{j}=0$ and exactly one $y_{ji}$ equals to $1$. In both cases, the second constraint of LP(11) is always satisfied. The third constraint is from the definition of makespan. By Remark 10, we know $p^{\prime}_{j}=p_{j}$ for every job. The last constraint indicates the total processing time of accepted jobs is bounded by $U$ and the jobs in $A^{\prime}_{1}$ has been accepted. Note that LP(11) has exactly $(m+1)|S|+m|T|\leq(m+1)n$ variables and $|T|+|S|+m+1\leq n+m+1$ constraints. The linear program is polynomial-time solvable in $n,m$ because the time complexity of a linear program is polynomial in the number of continuous variables and polynomial in the number of constraints [13].
We consider a basic solution of LP(11) and partition the tiny jobs into two sets as follows.
$T_{1}=\{J_{j}\in T:\mbox{ $y_{ji}=1$ for some $i$}\}\mbox{ and }T_{2}=T\setminus T_{1}.$ (12)
Similarly, we partition the safe jobs into three sets as follows.
$S_{1}=\{J_{j}\in S:\mbox{ $y_{ji}=1$ for some $i$}\},S_{2}=\{J_{j}\in S:\mbox{$x_{j}=1$}\},\mbox{ and }S_{3}=S\setminus(S_{1}\cup S_{2}).$ (13)
Lemma 20.
$|T_{2}|+|S_{3}|\leq m+1$.
Proof.
Since the number of constraints of LP(11) is $|T|+|S|+m+1$, there are at most $|T|+|S|+m+1$ positive values in the basic solution. By Eqs.(12) and (13), it is easy to see the number of positive values is at least $|T_{1}|+2|T_{2}|+|S_{1}|+|S_{2}|+2|S_{3}|=|T|+|S|+|T_{2}|+|S_{3}|$. Therefore, we have $|T|+|S|+|T_{2}|+|S_{3}|\leq|T|+|S|+m+1$ and thus the lemma is proved. ∎
We accept the jobs in $S_{1}$ and assign each job $J_{j}$ in $T_{1}\cup S_{1}$ to the machine $M_{i}$ such that $y_{ji}=1$. Finally, we reject all safe jobs in $S_{2}\cup S_{3}$ and for the tiny job set $T_{2}$, we arbitrarily append at most two of them at the end of the sub-schedule on each machine. By Lemma 20, $|T_{2}|\leq m+1$ and therefore, all jobs of $T_{2}$ have been assigned to the machines. We present the high-level description of the whole algorithm as follows.
Algorithm 2 Algorithm Approx2 for $Pm\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$
Input: An instance $J=\{J_{1},\ldots,J_{n}\}$, a fixed number $m$ of machines and a fixed constant $10<\epsilon<$ 1;
Call Approx1 in Algorithm 1 on the instance $I$ to obtain cost $Z\leq 2Z^{*}$;
use $Z$ to normalize instance $I$ so that $\frac{1}{2}\leq Z^{*}\leq Z=1$ (Eq.(8));
set up $\Delta$ using $\epsilon$ and $m$ as in Eq.(9) and partition jobs into risky and safe as in Definition 7;
set up $\delta$ using $\epsilon$ and $\Delta$ as in Eq.(10) and construct the restricted instance $I^{\prime}$as in Definition 9;
partition the risky jobs in $I^{\prime}$ into at most $\frac{1}{\delta}$ subsets $B_{i}$ as in Definition 12;
for each rejection type $(k_{1},\ldots,k_{1/\delta})$ as in Definition 13 do
     reject the first $k_{i}$ jobs in $B_{i}$ for every $i=1,\ldots,1/\delta$ and accept the other risky jobs;
     partition the accepted risky jobs into large and tiny as in Definition 16;
     for each large job assignment as in Definition 18 do
         assign each large job to the machine stated by the large job assignment;
         construct the linear program as in LP(11) and compute a basic solution;
         partition the tiny jobs into two sets $T_{1},T_{2}$ as in Eq.(12)
         partition the safe jobs into three sets $S_{1},S_{2},S_{3}$ as in Eq.(13);
         accept the jobs in $S_{1}$ and assign each job $J_{j}$ in $T_{1}\cup S_{1}$ to the machine $M_{i}$ such that $y_{ji}=1$;
         reject the jobs in $S_{2}\cup S_{3}$;
         append at most two jobs in $T_{2}$ at the end of the sub-schedule on each machine;
     end for
end for
return the feasible solution with the minimum cost for $I^{\prime}$;
replace back the original rejection penalty of each job in the selected solution.
Theorem 21.
Algorithm Approx2 is an EPTAS for $Pm\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$.
Proof.
We first prove the returned solution has a cost at most $(1+\epsilon)Z^{*}$. Let $\pi^{1}$ denote the schedule for $I^{\prime}$ when we consider the rejection type and the assignment of large jobs associated with the schedule $\pi^{\prime}$ stated in Lemma 15. By the linear program LP(11), the cost of $\pi^{1}$ is bounded by the cost of $\pi^{\prime}$ if we ignore the jobs in $T_{2}\cup S_{3}$. Note that each job $J_{j}\in S_{3}$ is safe and thus by Definition 7, $e_{j}\leq\Delta$. Since we reject the jobs in $S_{3}$, by Lemma 20, the increased rejection penalty is at most $(m+1)\Delta$. Moreover, each job $J_{j}\in T_{2}$ is tiny and thus by Definition 16, $p_{j}\leq\Delta$. Therefore, the increment on the makespan caused by the jobs in $T_{2}$ is at most $2\Delta$. By Lemma 15 and Eqs.(9), (8), the cost of $\pi^{1}$ is at most
$(1+\frac{\epsilon}{2})Z^{*}(I)+(m+3)\Delta\leq(1+\frac{\epsilon}{2})Z^{*}(I)+\frac{\epsilon}{4}\leq(1+\epsilon)Z^{*}(I).$
When we substitute back the original rejection penalties, the cost does not increase. Therefore, the cost of the solution produced is at most $(1+\epsilon)Z^{*}(I)$.
By Lemmas 14 and 19, the number of distinct rejection types and large job assignments are both fixed constants. The linear program is polynomial-time solvable in $m,n$. Therefore, Approx2 is an EPTAS for $Pm\mid rej,U,r_{i}=r\mid C_{\max}+\sum_{J_{j}\in R}e_{j}$. ∎
5 Conclusions
We study the scheduling with rejection problem where the total machine cost does not exceed a given upper bound $\hat{U}$. We assume the number $m$ of machines is a fixed constant and the problem is denoted as $Pm\mid rej,\hat{U},r_{i}=r\mid C_{\max}+\sum_{J_{j}\in R}e_{j}$. We present a simple $2$-approximation algorithm for $Pm\mid rej,\hat{U},r_{i}=r\mid C_{\max}+\sum_{J_{j}\in R}e_{j}$ and then further improve it to an EPTAS. Our EPTAS enumerates all possible rejection types and all possible large job assignment types to determine the rejection decisions for all risky jobs and the schedules of all accepted large risky jobs. Finally, we establish a linear program to deal with the safe and tiny jobs.
One possible future research direction is to design an FPTAS for $Pm\mid rej,\hat{U},r_{i}=r\mid C_{\max}+\sum_{J_{j}\in R}e_{j}$. Note that when $m$ is part of the input, the problem, denoted as $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}+\sum_{J_{j}\in R}e_{j}$, becomes strongly NP-hard. Therefore, another research direction is to extend our results to present a PTAS for $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}+\sum_{J_{j}\in R}e_{j}$. Moreover, we currently assume the energy cost rates are uniform and it is natural to think about the non-uniform energy cost rates.
Acknowledgments
This work has received funding from NSF award 2309902.